OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 4 and 9. What is the length of the altitude? 9 13 4 6

OpenStudy (anonymous):

 So, all we have to do is to find x for this question. All you need is Pyth. Thm. here (Of course with some additions.) In triangle ABC, a^2 + b^2 = (4+9)^2 a^2 + b^2 = 13^2 ---(1) In triangle ACD, x^2 + 4^2 = b^2 ---(2) In triangle BCD, x^2 + 9^2 = a^2 ---(3) (2) + (3), (x^2 + 4^2) + (x^2 + 9^2) = a^2 + b^2 So, you can sub (1) to the right side to solve x.